• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 00:24:19 on January13,2002

    Entry number 80161

    keyword=cryotarget cool-down

    Dave Meekins, Chris Keith, Mike Seely and Sue Witherspoon worked during the
    day to fix problems with the target control system. With a few problems remain
    (to be fixed later), Mike Seely started cool-down. I came in at around 6:30 pm.
    Took over at about 7:00pm. At the time, there remain several problems, which
    I will list at the end. Target was at below 30 K already, but liquification not started.
    I continued the cooldown with assistance from Louis Bimbot. With cooling line temperature sensors (diodes) not working, I had to judge from other indications
    to decide when to switch to cold return. With target at about 24 K for some time,
    the switch from warm return to cold return were completed at about 8:40 pm.
    Target was full at about 10:30 pm and was only at 15-16 psi at 19 K. So we added
    hydrogen from the bottle to about 25 psi. And adjusted target fan to 60 Hz. Target
    cooldown completed at 11:00 pm. Here are the remain problems*:
    1*) JT-valve computer control not working, Dave is currently working on that.
    *(Dave has temperary fixed it*).
    2*) Problem with startup script for the alarm and chart. Charts can still be accessed
    through the main control gui. Dave is also working on the alarm startup. And will
    also set the alarm values.
    3*) Lifter software has problems. None of the position buttens are working. The only
    way to move the target is to type in the encoder number (Mike wanted ONLY
    the target 'expert' to do the target moving).
    4*) Loop fan computer control is not working. It is locally set to 60 Hz and we
    probably do not need to change it unless problem happens.
    5*) Diode temeprature sensors for cooling lines are not working.
    6*) High power heater mode indicator is not working. For loop2, the PID indicator always on.
    7*) Chart display vertical axis change not working.
    8*) Default value values for heater PID is lost *(loaded to be 0*), need to either find the old values or play with it to get the optimal values.
    9*) Cool-down motion was not monitored.
    10*) No camaras for the instrumentation panels.
    The target group will continue work on fixing the problems later.